OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. sqrt(s^8)+sqrt(25s^8)+2sqrt(s^8)-sqrt(s^4). It's multiple choice.

OpenStudy (anonymous):

\[\sqrt{s ^{8}} +\sqrt{25s ^{8}} +2\sqrt{s ^{8}} - \sqrt{s ^{4}} ?\]